# OCCAM: Towards Cost-Efficient and Accuracy-Aware Image Classification Inference

### v1

@@ Abstract
Image classification is a fundamental building block for a majority of computer vision applications. With the growing popularity and capacity of machine learning models, people can easily access trained image classifiers as a service online or offline. However, model use comes with a cost and classifiers of higher capacity usually incur higher inference costs. To harness the respective strengths of different classifiers, we propose a principled approach, OCCAM, to compute the best classifier assignment strategy over image classification queries (termed as the optimal model portfolio) so that the aggregated accuracy is maximized, under user-specified cost budgets. Our approach uses an unbiased and low-variance accuracy estimator and effectively computes the optimal solution by solving an integer linear programming problem. On a variety of real-world datasets, OCCAM achieves $40\%$ cost reduction with little to no accuracy drop.
@@ 1 Introduction
With the breakthroughs in AI and advances in computer hardware (e.g., GPUs and TPUs) in recent decades, applications of computer vision have permeated our daily lives, ranging from face recognition systems to autonomous driving technologies. Among all the day-to-day computer vision applications, a fundamental building block is the task of image classification, where given an image, an algorithm needs to recognize the object content inside the image.
The task of image classification has a long history in the computer vision literature. Before the emergence of deep learning, people mainly focused on designing handcrafted features or descriptors for images, such as HOG [Dalal and Triggs, 2005] and SIFT [Lowe, 2004]. With the growing capability of deep learning models, many neural network architectures including convolutional neural networks (CNNs) [LeCun et al., 1998] and Transformers [Vaswani et al., 2017] have been proposed, e.g., AlexNet [Krizhevsky et al., 2012], ResNet [He et al., 2016], Vision Transformer [Dosovitskiy et al., 2020], and Swin Transformer [Liu et al., 2021]. Though larger neural network models are equipped with higher capacity, they often come with higher costs as well, e.g., hardware usage and latency (time), for both training and inference. This can potentially impose an enormous cost on both end users of image classification services and the service providers (e.g., Google [^1] https://cloud.google.com/prediction, Amazon [^2] https://aws.amazon.com/machine-learning, and Microsoft [^3] https://studio.azureml.net). In response to this challenge, there has been a notable surge in interest directed towards the development of smaller, cost-effective image classifiers, e.g., MobileNet [Howard et al., 2017], where depthwise separable convolutions are used to trade classification accuracy for efficiency. However, empirical evaluations conducted in [Su et al., 2018], as well as our own independent assessment (see Figure 1(a)), consistently indicate that smaller models tend to exhibit a gap in classification accuracy compared to their larger counterparts.
@@
Confronted with the general tradeoff between classification accuracy and inference cost, we advocate a hybrid inference framework which seeks to combine the advantages of both small and large models. Specifically, we study the problem, given a user specified cost budget and a group of ML classifiers of different capacity and cost, assign classifiers to resolve different image classification queries so that the aggregated accuracy is maximized and the overall cost is under the budget. We formally define it as the optimal model portfolio problem (details in Section 3). Our approach is motivated by the observation that while small classifiers typically have reduced accuracy over the population, they can still agree with large classifiers on certain queries a large proportion of the time, which suggests the existence of a subset of “easy” queries on which even small classifiers can make the right prediction. This is also illustrated in Figure 1(b) where we plot the frequency with which different classifiers successfully make the right prediction on the same image queries. For instance, ResNet-18 [He et al., 2016] can correctly classify $75\%$ of the images on which SwinV2-B [Liu et al., 2022] makes the right prediction, suggesting that we can replace SwinV2-B with ResNet-18 on these image queries, saving significant inference costs without any accuracy drop (details in Section 5).
@@
With this insight, we propose a principled approach, Optimization with Cost Constraints for Accuracy Maximization (OCCAM), to effectively identify easy queries and assign classifiers to different user queries to maximize the overall classification accuracy subject to the given cost budgets. We present an unbiased and low-variance estimator for classifier test accuracy with asymptotic guarantees. The intuition is that for well-separated classification problems such as image classification [Yang et al., 2020], we can learn robust classifiers that have similar performance on similar queries. For each query image, we compute its nearest neighbours in pre-computed samples to estimate the test accuracy for each classifier. Previous work [Chen et al., 2022] trains ML models to predict the accuracy, which requires sophisticated configuration and lacks performance guarantees that are critical in real-world scenarios. To our best knowledge, we are the first to open up the black box by developing a white-box accuracy estimator for ML classifiers with statistical guarantees. Next, armed with our classifier accuracy estimator, we compute the optimal classifier assignment strategy over all query images (optimal model portfolio) subject to a given cost budget by solving an integer linear programming (ILP) problem (see Section 4). As a preview, Figure 1(c) shows that OCCAM can achieve $20\%$ cost reduction with less than $1\%$ accuracy drop. We show even higher cost reduction with little to no accuracy drop on various real-world datasets in Section 5. Figure 2 depicts the overall pipeline of OCCAM.
@@
Our main technical contributions are: (1) we formally define the optimal model portfolio problem to reduce overall inference costs while maintaining high performance subject to user-specified cost budgets (Section 3); (2) we propose a novel and principled approach, OCCAM, to effectively compute the optimal model portfolio with statistical guarantees (Section 4); and (3) we provide an extensive experimental evaluation on a variety of real-world datasets on the image classification task (Section 5) demonstrating the effectiveness of OCCAM.
@@ 2 Related Work
Image Classification. Image classification is a fundamental task in computer vision, where given an image, a label needs to be predicted. It serves as an essential building block for many high-level AI tasks, e.g., image captioning [Vinyals et al., 2015] and visual question answering [Antol et al., 2015], where objects need to be first recognized. Before the deep learning era, researchers mainly adopted statistical methods with handcrafted features for the task, e.g., SIFT [Lowe, 2004]. With the growing capacity of deep learning models, from convolutional neural networks (CNN) [Krizhevsky et al., 2012, Simonyan and Zisserman, 2014, He et al., 2016, Szegedy et al., 2016] to Transformer architectures [Dosovitskiy et al., 2020, Liu et al., 2021], the classification accuracy on standard image classification benchmarks [Krizhevsky et al., 2009, Russakovsky et al., 2015] has been greatly improved. In this work, we utilize both CNN (e.g., ResNet models) and Transformer (e.g., Swin Transformers) image classifiers to illustrate and evaluate our proposed approach, OCCAM.
@@
Efficient Machine Learning (ML) Inference. Efficient ML inference is crucial for real-time decision-making in various applications such as autonomous vehicles [Tang et al., 2021], healthcare [Miotto et al., 2018], and fraud detection [Alghofaili et al., 2020]. It involves applying a pre-trained ML model to make predictions, where the inference cost is expected to dominate the overall cost incurred by the model. Model compression, which replaces a large model with a smaller model of comparable accuracy, is the most common approach employed for enhancing ML inference efficiency. Common techniques for model compression include model pruning [Hassibi et al., 1993, LeCun et al., 1989], quantization [Jacob et al., 2018, Vanhoucke et al., 2011], knowledge distillation [Hinton et al., 2015, Urban et al., 2016], neural architecture search [Elsken et al., 2019, Zoph and Le, 2016], and so on. These static efficiency optimizations typically lead to a fixed model with lower inference cost but also reduced accuracy compared to its larger counterpart, which may not suffice in highly sensitive applications like collision detection [Wang et al., 2021] and prognosis prediction [Zhu et al., 2020]. This shortcoming is already evident in the inference platforms discussed in Section 1, highlighting the need for more dynamic optimizations to effectively address the diverse demands of users.
@@
Hybrid ML Inference. Recent works [Kag et al., 2022, Ding et al., 2022, 2024] have introduced a novel inference paradigm termed hybrid inference, which invokes models of different sizes on different queries, as opposed to employing a single model on all inference queries. The smaller model generally incurs a lower inference cost but also exhibits reduced accuracy compared to the larger model. The key idea is to identify easy inference queries on which the small models are likely to make correct predictions and invoke small models on them when cost budgets are limited, thereby reducing overall inference costs while preserving solution accuracy. By adjusting the cost budgets, users can dynamically trade off between accuracy and cost within the same inference setup.  [Kag et al., 2022, Ding et al., 2022, 2024] consider a simple setting of only one large and one small model and do not allow for explicit cost budget specification, which could be necessary for production scenarios. [Chen et al., 2020] studies a setup with multiple ML models and learns an adaptive strategy to generate predictions by calling a base model and sometimes an add-on model when the base model quality scores are lower than the learned thresholds. However, both base and add-on models are selected in a probabilistic manner and this approach fails to satisfy the user-specified cost budgets deterministically. [Chen et al., 2022] studies a similar setup with multiple ML models and allocates cost budgets according to model-based accuracy prediction. This approach requires a separate training phase for the accuracy predictor, which needs a large amount of training data, and provides no guarantee on the prediction quality. Unlike previous works, we propose an unbiased and low-variance accuracy estimator with asymptotic guarantees, based on which we present a novel approach, OCCAM, to effectively compute the optimal assignment of classifiers to given queries, under the cost budgets given by users.
@@ 4.1 Estimating $Accuracy(\mu)$
Previous work on Hybrid ML [Kag et al., 2022, Ding et al., 2022, 2024, Chen et al., 2022] typically relies on training a neural router to predict the accuracy of a given set of classifiers for given user queries, based on which queries are routed to different classifiers. Such a paradigm not only involves a non-trivial training configuration but also lacks estimation guarantees which can be critical in scientific and production settings. We propose a principled approach to estimate the test accuracy of a given model portfolio for given user queries. By leveraging the specific structure of well-separated classification problems like image classification, we propose an unbiased low-variance estimator for the test accuracy with asymptotic guarantees.
@@ 4.2 Computing $\mu^{*}$ with $Accuracy(\mu)$
With the estimator in place, we formulate the problem of finding the optimal model portfolio as an integer linear programming (ILP) problem as follows. Given a set of $M$ classifiers $f_{1},f_{2},\cdots,f_{M}$, user queries $X=\{x_{1},x_{2},\cdots,x_{N}\}$, pre-computed samples $S_{1},S_{2},\cdots,S_{K}$, and budget $B\in\mathbb{R}^{+}$, we have the following ILP problem  [^6] Our problem can be rephrased as “selecting for each query image, one item (i.e., ML classifier) from a collection (the set of all classifiers) so as to maximize the total value (accuracy) while adhering to a predefined weight limit (cost budget)”, which is a classic multiple choice knapsack problem (MCKP) [Kellerer et al., 2004] and the ILP formulation is the natural choice..
$\begin{split}\max\quad&\sum_{i=1}^{M}\sum_{j=1}^{N}\widehat{SP}_{i}(x_{j})\cdot t_{i,j}\\ \textrm{s.t.}\quad&\sum_{i=1}^{M}\sum_{j=1}^{N}b_{i}\cdot t_{i,j}\leq B\\ &\sum_{i=1}^{M}t_{i,j}=1\text{ for }j=1,2,\cdots,N,\text{ and }t_{i,j}\in\{0,1\}\\ \end{split}$ (4)
where $t_{i,j}$ are boolean variables and $t_{i,j}=1$ iff the classifier $f_{i}$ is assigned to query $x_{j}$. Clearly, the optimal model portfolio $\mu^{*}$ can be efficiently computed as $\mu^{*}(x_{j})=i$ iff $t^{*}_{i,j}=1$, for $i\in[M]$ and $j\in[N]$, where $t^{*}_{i,j}$ is the optimal solution to the ILP problem above. While ILP problems are NP-hard in general, we can use standard ILP solvers (e.g., HiGHS [Huangfu and Hall, 2018]) to efficiently compute the optimal solution in practice.
@@ 5.1 Evaluation Setup
Task. We consider the image classification task: given an image, predict a class label from a set of predefined class categories. We assume that each image has a unique ground-truth class label.
Datasets. We consider 4 widely studied datasets for image classification: CIFAR-10 (10 classes) [Krizhevsky et al., 2009], CIFAR-100 (100 classes) [Krizhevsky et al., 2009], Tiny ImageNet (200 classes) [CS231n, ], and ImageNet-1K (1000 classes) [Russakovsky et al., 2015]. Both CIFAR-10 and CIFAR-100 contain $50,000$ training images and $10,000$ test images. Tiny ImageNet contains $100,000$ training images and $10,000$ validation images, and ImageNet-1K has 1,281,167 training images and 50,000 validation images. We use the test splits of CIFAR-10 and CIFAR-100 as well as the validation splits of Tiny ImageNet and ImageNet-1K for evaluation purposes. Details of those datasets are in LABEL:{sec:app_image_dset}.
Models. We consider a total of 7 classifiers: ResNet-[18, 34, 50, 101] [He et al., 2016] [^7] Numbers in bracket indicate the model’s layer number. and SwinV2-[T, S, B] [Liu et al., 2022] [^8] Letters in bracket indicate the Swin Transformer V2’s size. T/S/B means tiny/small/base. Among these classifiers, ResNet-18 is the smallest (in terms of number of model parameters and training/inference time) and thus has the least capacity, while SwinV2-B is the largest and with the highest accuracy in general (see Figure 1(a)). We take the classifiers pre-trained on the ImageNet-1K dataset [Russakovsky et al., 2015]. We directly use the pre-trained models on ImageNet-1K, while on other datasets, we freeze everything but train only the last layer from scratch. The output dimension of the last layer is set to be the same as the number of image classes on the test dataset. We implement the soft classifier (see Section 4.1) by sampling w.r.t. the probability distribution output by the softmax layer, i.e., $Pr[f_{i}(x)=j]=\frac{exp(z_{j}/\tau)}{\sum_{k}exp(z_{k}/\tau)}$, where $z_{k}$ is the logit for $k\in[C]$ and $\tau$ is the hyper-parameter temperature controlling the randomness of predictions. We choose a small $\tau$ (1e-3) to reduce the variance in predictions. At test time, to obtain consistent results, all classifiers make predictions by outputting the most likely class labels (i.e., $\arg\max_{j}f_{i}(x)[j]$), equivalently to having soft classifiers with $\tau\to 0$. Model training details are in Section C.2. All experiments are conducted with one NVIDIA V100 GPU of 32GB GPU RAM. Codes will be released upon acceptance.
@@
Baselines. We compare our approach with three baselines: single best, random, and FrugalMCT [Chen et al., 2022]. Single best always chooses the strongest (i.e., most expensive) model for a given cost budget. Random estimates classifier accuracy with random guesses (i.e., uniform samples from $[0,1]$) and solves the problem in Equation 4 with the same ILP solver as ours. FrugalMCT [Chen et al., 2022] is a recent work which selects the best ML models for given user budgets in an online setting, using model-based accuracy estimation. Following the same setting in [Chen et al., 2022], we train random forest regressors on top of the model-extracted features (e.g., ResNet-18 features), as the accuracy predictor. The predicted accuracy is used in Equation 4, which is solved by the same ILP solver as ours.
Our Method. We evaluate OCCAM (see Section 4) under various metrics (i.e., $l_{1}$, $l_{2}$, and $l_{\infty}$ norms) and cost budgets. We consider images represented by model-based embeddings. Specifically, we extract the image feature [^9] The image feature is the last layer output of a ML model (e.g., ResNet-18) trained on the target dataset, given an input image. of the query image and all the validation images. The costs incurred by feature extraction are deducted from the user budget $B$ before we compute the optimal model portfolio. We report the test accuracy under different cost budgets for OCCAM and all baselines in Section 5.2 (Figure 3 and Table 2), validate that OCCAM is cost-aware and indeed selecting the most profitable ML models to deliver high accuracy solutions in Section 5.3 (Figure 4(a)), demonstrate the effectiveness of OCCAM with limited samples in Section 5.4 (Figure 4(b)), investigate the nearest neighbour distance with different sample sizes in Section A.2, show that the estimation error of our accuracy estimator quickly decreases as the sample size increases in Section A.3, test the generalizability of OCCAM with different feature extractors in Section A.4, and provide more performance results under different metrics in Section A.5.

### v2

# OCCAM: Towards Cost-Efficient and Accuracy-Aware Classification Inference

@@ Abstract
Classification tasks play a fundamental role in various applications, spanning domains such as healthcare, natural language processing and computer vision. With the growing popularity and capacity of machine learning models, people can easily access trained classifiers as a service online or offline. However, model use comes with a cost and classifiers of higher capacity (such as large foundation models) usually incur higher inference costs. To harness the respective strengths of different classifiers, we propose a principled approach, OCCAM, to compute the best classifier assignment strategy over classification queries (termed as the optimal model portfolio) so that the aggregated accuracy is maximized, under user-specified cost budgets. Our approach uses an unbiased and low-variance accuracy estimator and effectively computes the optimal solution by solving an integer linear programming problem. On a variety of real-world datasets, OCCAM achieves $40\%$ cost reduction with little to no accuracy drop.
@@ 1 Introduction
Classification is a fundamental task with numerous real-world applications, from medical diagnosis [Abhisheka et al., 2024] to natural language processing [Kowsari et al., 2019] and object recognition [Zou et al., 2023]. In recent years, a wealth of pre-trained classifiers has become accessible, both online and offline, at varying costs. For instance, Google’s Facial Detection service [Google, ] charges \$1.50 per 1,000 images, providing developers with a reliable solution for face recognition. Simultaneously, a variety of open-source models, such as ResNet [He et al., 2016], Vision Transformer [Dosovitskiy et al., 2020], and Swin Transformer [Liu et al., 2021], have been made available for developers to integrate into their services.
On the one hand, empirical evaluations conducted in [Su et al., 2018] and our independent assessment (see Figure 1(a)), consistently indicate that smaller/cheaper models tend to exhibit a gap in classification accuracy compared to their larger counterparts. On the other hand, though larger neural models are equipped with higher capacity, they often come with higher costs, e.g., hardware usage and latency, for both training and inference. This can potentially impose an enormous cost on both end users of classification services and the service providers (e.g., Google [^1] https://cloud.google.com/prediction, Amazon [^2] https://aws.amazon.com/machine-learning, and Microsoft [^3] https://studio.azureml.net).
Confronted with the general trade-off between classification accuracy and inference cost, we advocate a hybrid inference framework which seeks to combine the advantages of both small and large models. Specifically, we study the problem, how to optimally select and assign models to different classification queries in order to achieve the highest accuracy while adhering to a limited cost budget? We formally refer to it as the optimal model portfolio problem (details in Section 3). Our approach is motivated by the observation that while small classifiers typically have reduced accuracy over the population, they can still agree with large classifiers on certain queries a large proportion of the time, which suggests the existence of a subset of “easy” queries on which even small classifiers can make the right prediction. This is also illustrated in Figure 1(b) where we plot the frequency with which different classifiers successfully make the right prediction on the same classification queries. Taking image classification as an example, a small classifier, ResNet-18 [He et al., 2016], can correctly classify $75\%$ of the images on which the large classifier, SwinV2-B [Liu et al., 2022], makes the right prediction, suggesting that we can replace SwinV2-B with ResNet-18 on these image queries, saving significant inference costs without any accuracy drop (details in Section 5).
@@
With this insight, we propose a principled approach, Optimization with Cost Constraints for Accuracy Maximization (OCCAM), to effectively identify easy queries and assign classifiers to different user queries to maximize the overall classification accuracy subject to the given cost budgets. Previous work [Chen et al., 2022] trains ML models to select classifiers, which requires sophisticated configuration and lacks performance guarantees that are critical in real-world scenarios. In this paper, by leveraging weak assumptions, such as the well-separation assumption [Yang et al., 2020], we demonstrate that it is possible to compute optimal model assignments with statistical guarantees. As an example, in previous work [Yang et al., 2020], the image classification task [Singh and Singh, 2020], a critical and well-studied application of classification, has been observed to empirically satisfy the well-separation assumption, which is also corroborated by our own independent evaluation (see Section A.1). The intuition is that for well-separated classification problems such as image classification, we can learn robust classifiers that have similar performance on similar queries. By leveraging the specific structure of well-separated classification tasks, we develop an unbiased and low-variance estimator for classifier test accuracy with asymptotic guarantees. For each classification query, we compute its nearest neighbours in pre-computed samples to estimate the test accuracy for each classifier. To our best knowledge, we are the first to open up the black box by developing a white-box accuracy estimator for ML classifiers with statistical guarantees. Next, armed with our classifier accuracy estimator, we compute the optimal classifier assignment strategy over all classification queries (optimal model portfolio) subject to a given cost budget by solving an integer linear programming (ILP) problem (see Section 4). As a preview, Figure 1(c) shows that OCCAM can achieve $20\%$ cost reduction with less than $1\%$ accuracy drop. We show even higher cost reduction with little to no accuracy drop on various real-world datasets in Section 5. Figure 2 depicts the overall pipeline of OCCAM: (1) draw samples and pre-compute the ML classifier accuracy for each sample, (2) for each test query, use its nearest neighbour from all samples to estimate the test accuracy of each classifier, (3) compute the optimal model portfolio w.r.t. cost budgets by solving the ILP problem.
@@
Our main technical contributions are: (1) we formally define the optimal model portfolio problem to reduce overall inference costs while maintaining high performance subject to user-specified cost budgets (Section 3); (2) we propose a novel and principled approach, OCCAM, to effectively compute the optimal model portfolio with statistical guarantees under weak assumptions (Section 4); and (3) as an illustration of the usefulness of OCCAM, we conduct extensive experiments on the image classification task given its well-separation property and demonstrate the effectiveness of OCCAM on a variety of real-world datasets (Section 5).
@@ 2 Related Work
Efficient Machine Learning (ML) Inference. Efficient ML inference is crucial for real-time decision-making in various applications such as autonomous vehicles [Tang et al., 2021], healthcare [Miotto et al., 2018], and fraud detection [Alghofaili et al., 2020]. It involves applying a pre-trained ML model to make predictions, where the inference cost is expected to dominate the overall cost incurred by the model. Model compression, which replaces a large model with a smaller model of comparable accuracy, is the most common approach employed for enhancing ML inference efficiency. Common techniques for model compression include model pruning [Hassibi et al., 1993, LeCun et al., 1989, Ding et al., ], quantization [Jacob et al., 2018, Vanhoucke et al., 2011], knowledge distillation [Hinton et al., 2015, Urban et al., 2016], and neural architecture search [Elsken et al., 2019, Zoph and Le, 2016]. These static efficiency optimizations typically lead to a fixed model with lower inference cost but also reduced accuracy compared to its larger counterpart, which may not suffice in highly sensitive applications like collision detection [Wang et al., 2021] and prognosis prediction [Zhu et al., 2020]. This shortcoming is already evident in the inference platforms discussed in Section 1, highlighting the need for more dynamic optimizations to effectively address the diverse demands of users.
@@
Hybrid ML Inference. Recent works [Kag et al., 2022, Ding et al., 2022, 2024] have introduced a novel inference paradigm termed hybrid inference, which invokes models of different sizes on different queries, as opposed to employing a single model on all inference queries. The smaller model generally incurs a lower inference cost but also exhibits reduced accuracy compared to the larger model. The key idea is to identify easy inference queries on which the small models are likely to make correct predictions and invoke small models on them when cost budgets are limited, thereby reducing overall inference costs while preserving solution accuracy. By adjusting the cost budgets, users can dynamically trade off between accuracy and cost within the same inference setup.  Kag et al. [2022], Ding et al. [2022, 2024] consider a simple setting of only one large and one small model and do not allow for explicit cost budget specification, which could be necessary for production scenarios. Chen et al. [2020] studies a setup with multiple ML models and learns an adaptive strategy to generate predictions by calling a base model and sometimes an add-on model when the base model quality scores are lower than the learned thresholds. However, both base and add-on models are selected in a probabilistic manner and this approach fails to satisfy the user-specified cost budgets deterministically. Chen et al. [2022] studies a similar setup with multiple ML models and allocates cost budgets according to model-based accuracy prediction. This approach requires a separate training phase for the accuracy predictor, which needs a large amount of training data, and provides no guarantee on the prediction quality. Mixture-of-experts (MoE) [Shazeer et al., 2017] also involves dynamically deciding which model (or “expert”) processes a given query based on specific criteria. However, MoE often assumes homogenous experts [Fedus et al., 2022] with routing decisions made by learned gating functions that require significant computational resources to re-train if we want to add/update/delete experts. Unlike previous works, we propose an unbiased and low-variance accuracy estimator with asymptotic guarantees, based on which we present a novel training-free approach, OCCAM, to effectively compute the optimal assignment of diverse classifiers to given queries, under the cost budgets specified by users. It is worth pointing out that OCCAM is a an instance from the Algorithm Selection problem [Rice, 1976] with a specific focus on classification tasks and pre-trained classifiers.
Image Classification. Image classification is a fundamental task in computer vision, where given an image, a label needs to be predicted. It serves as an essential building block for many high-level AI tasks, e.g., image captioning [Vinyals et al., 2015] and visual question answering [Antol et al., 2015], where objects need to be first recognized. With the growing capacity of deep learning models, from convolutional neural networks (CNN) [Krizhevsky et al., 2012, Simonyan and Zisserman, 2014] to Transformer architectures [Dosovitskiy et al., 2020, Liu et al., 2021], the classification accuracy on standard image classification benchmarks [Russakovsky et al., 2015] has been greatly improved. In this work, we use the image classification task to illustrate and evaluate our proposed approach OCCAM given its well-separation property [Yang et al., 2020]. We utilize both CNN (e.g., ResNet models) and Transformer (e.g., Swin Transformers) image classifiers in our evaluation.
@@ 4.1 Estimating $Accuracy(\mu)$
Previous work on Hybrid ML [Kag et al., 2022, Chen et al., 2022, Ding et al., 2024] typically relies on training a neural router to predict the accuracy of a given set of classifiers for given user queries, based on which queries are routed to different classifiers. Such a paradigm not only involves a non-trivial training configuration but also lacks estimation guarantees which can be critical in scientific and production settings. We propose a principled approach to estimate the test accuracy of a given model portfolio for given user queries. By leveraging the specific structure of well-separated classification problems like image classification, we propose an unbiased low-variance estimator for the test accuracy with asymptotic guarantees.
@@ 4.2 Computing $\mu^{*}$ with $Accuracy(\mu)$
With the estimator in place, we formulate the problem of finding the optimal model portfolio as an integer linear programming (ILP) problem as follows. Given a set of $M$ classifiers $f_{1},f_{2},\cdots,f_{M}$, user queries $X=\{x_{1},x_{2},\cdots,x_{N}\}$, pre-computed samples $S_{1},S_{2},\cdots,S_{K}$, and budget $B\in\mathbb{R}^{+}$, we have the following ILP problem  [^6] Our problem can be rephrased as “selecting for each query, one item (i.e., ML classifier) from a collection (the set of all classifiers) so as to maximize the total value (accuracy) while adhering to a predefined weight limit (cost budget)”, which is a classic multiple choice knapsack problem (MCKP) [Kellerer et al., 2004] and the ILP formulation is the natural choice..
$\max\,\sum_{i=1}^{M}\sum_{j=1}^{N}\widehat{SP}_{i}(x_{j})\cdot t_{i,j}\quad\textrm{s.t.}\,\sum_{i=1}^{M}\sum_{j=1}^{N}b_{i}\cdot t_{i,j}\leq B\;\text{and}\,\sum_{i=1}^{M}t_{i,j}=1,\text{ for }j=1,2,\cdots,N$ (4)
where $t_{i,j}\in\{0,1\}$ are boolean variables and $t_{i,j}=1$ iff the classifier $f_{i}$ is assigned to query $x_{j}$. Clearly, the optimal model portfolio $\mu^{*}$ can be efficiently computed as $\mu^{*}(x_{j})=i$ iff $t^{*}_{i,j}=1$, for $i\in[M]$ and $j\in[N]$, where $t^{*}_{i,j}$ is the optimal solution to the ILP problem above. While ILP problems are NP-hard in general, we can use standard ILP solvers (e.g., HiGHS [Huangfu and Hall, 2018]) to efficiently compute the optimal solution in practice.
@@
The overall algorithm is shown in Algorithm 1. We first pre-compute the ML classifier accuracy for all samples (line 1-3). Next, for each test query, we use its nearest neighbour from all samples to estimate the test accuracy of each classifier (line 4-6). In line 7, we compute the optimal model portfolio with the accuracy estimation subject to cost budgets by solving the ILP problem (see Equation 4).
Input: test query batch $X$; ML classifiers $f_{1},f_{2},\cdots,f_{M}$ and costs $c_{1},c_{2},\cdots,c_{M}$; query samples $S_{1},S_{2},…,S_{k}$; user cost budget $B$.
Output: optimal model portfolio $\mu^{*}:X\to[M]$.
1
// Pre-compute the ML classifier accuracy for all samples
2 for $s\in S_{1}\cup S_{2}\cup\cdots\cup S_{k}$ do
3       for $i\in 1,2,\cdots,M$ do
4             $SampleAccuracy[s,i]=Accuracy(f_{i}(s),ground\_truth(s))$
5
6
// Use samples to estimate the accuracy for each test query
7 for $x\in X$ do
8       for $i\in 1,2,\cdots,M$ do
9             $AccuracyEstimation[x,i]=Mean(\{SampleAccuracy[NearestNeighbour(x,S_{j}),i],\text{for }1\leq j\leq k\})$
10
11
// Compute the optimal model portfolio with the accuracy estimation subject to the cost budget (see Equation 4)
12 $\mu^{*}=ILPSolver(AccuracyEstimation,c_{1},c_{2},\cdots,c_{M},B)$
return $\mu^{*}$
Algorithm 1 OCCAM Algorithm.
@@ 5.1 Evaluation Setup
Task. We consider the image classification task provided with its well-separation property [Yang et al., 2020]: given an image, predict a class label from a set of predefined class categories. We assume that each image has a unique ground-truth class label.
Datasets. We consider 4 widely studied datasets for image classification: CIFAR-10 (10 classes) [Krizhevsky et al., 2009], CIFAR-100 (100 classes) [Krizhevsky et al., 2009], Tiny ImageNet (200 classes) [CS231n, ], and ImageNet-1K (1000 classes) [Russakovsky et al., 2015]. Details of those datasets are in LABEL:{sec:app_image_dset}.
Models. We consider a total of 7 classifiers: ResNet-[18, 34, 50, 101] [He et al., 2016] [^7] Numbers in bracket indicate the model’s layer number. and SwinV2-[T, S, B] [Liu et al., 2022] [^8] Letters in bracket indicate the Swin Transformer V2’s size. T/S/B means tiny/small/base. Among these classifiers, ResNet-18 is the smallest (in terms of number of model parameters and training/inference time) and thus has the least capacity, while SwinV2-B is the largest and with the highest accuracy in general (see Figure 1(a)). We take the classifiers pre-trained on the ImageNet-1K dataset [Russakovsky et al., 2015]. We directly use the pre-trained models on ImageNet-1K, while on other datasets, we freeze everything but train only the last layer from scratch. Model training details are in Section C.2. All experiments are conducted with one NVIDIA V100 GPU of 32GB GPU RAM. Codes are available in https://github.com/DujianDing/OCCAM.git.
@@
Baselines. We compare our approach with three baselines: single best, random, and FrugalMCT [Chen et al., 2022]. Single best always chooses the strongest (i.e., most expensive) model for a given cost budget. Random estimates classifier accuracy with random guesses (i.e., uniform samples from $[0,1]$) and solves the problem in Equation 4 with the same ILP solver as ours. FrugalMCT [Chen et al., 2022] is a recent work which selects the best ML models for given user budgets in an online setting, using model-based accuracy estimation. Following the same setting in [Chen et al., 2022], we train random forest regressors on top of the model-extracted features (e.g., ResNet-18 features), as the accuracy predictor. The predicted accuracy is used in Equation 4, which is solved by the same ILP solver as ours to make a fair comparison in terms of solution quality and costs.
Our Method. We evaluate OCCAM (see Section 4) under various metrics (i.e., $l_{1}$, $l_{2}$, and $l_{\infty}$ norms) and cost budgets. We consider images represented by model-based embeddings. Specifically, we extract the image feature  [^9] The image feature is the last layer output of a ML model (e.g., ResNet-18) trained on the target dataset, given an input image. of the query image and all the validation images. The costs incurred by feature extraction are deducted from the user budget $B$ before we compute the optimal model portfolio. We report the test accuracy under different cost budgets for OCCAM and all baselines in Section 5.2 (Figure 3 and Table 2), validate that OCCAM is cost-aware and indeed selecting the most profitable ML models to deliver high accuracy solutions in Section 5.3 (Figure 4(a)), demonstrate the effectiveness of OCCAM with limited samples and various $\lambda$ values in Section 5.4 (Figures 4(b) and 4(c)), investigate the nearest neighbour distance in Section A.2, show that the estimation error of our accuracy estimator quickly decreases as the sample size increases in Section A.3, test the generalizability of OCCAM with different feature extractors in Section A.4, provide more performance results under different metrics (see Section A.5) and more classifiers (see Section A.6), show how to effectively choose $K$ and $s$ in Section A.7, perform overhead and upper bound performance analysis of OCCAM in Sections A.8 and A.9, and examine the classifier complementarity in Section A.10.